OpenStudy (anonymous):

Is x^2-4 a prime polynomial?

jimthompson5910 (jim_thompson5910):

Hint: a^2 - b^2 = (a-b)(a+b) by the difference of squares rule

OpenStudy (anonymous):

So its not?

OpenStudy (anonymous):

Oh wait it is ?

jimthompson5910 (jim_thompson5910):

can you write x^2-4 in the form a^2 - b^2

OpenStudy (anonymous):

no you cant

jimthompson5910 (jim_thompson5910):

are you sure?

OpenStudy (anonymous):

you cant make -4 into -b^2 unless you turn it to 2^2 ?

OpenStudy (anonymous):

-2^2

jimthompson5910 (jim_thompson5910):

4 = 2^2 So x^2 - 4 becomes x^2 - 2^2

jimthompson5910 (jim_thompson5910):

so we actually can rewrite x^2 - 4 into a^2 - b^2 form

jimthompson5910 (jim_thompson5910):

where in this case, a = x and b = 2